OpenStudy (anonymous):

PLEASE HELP I NEED DESPERATELY HOW TO DO THESE PROBLEMS I NEED TO ANSWER LIKE 30 OF THEM ON A TEST How do I solve the system of inequalities by graphing? for example a problem like this y less than or equal to x – 2 –4x + y more than sign – 3

OpenStudy (anonymous):

try using this website it's simple and it helped me with graphs http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html

OpenStudy (anonymous):

can you help me solve it because I don't understand how we are suppose to do it

OpenStudy (anonymous):

does it do with graphing calculators at all?

OpenStudy (anonymous):

hang on il send a picture

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

http://my.hrw.com/math06_07/nsmedia/tools/Graph_Calculator/graphCalc.html I don't know if this will show you but you want to plug it in like I have it.

OpenStudy (anonymous):

The top number you want to plug in the x -2 and the bottom number should be -4xy - 3 Like i said the website will help you out

OpenStudy (anonymous):

Im confused

OpenStudy (anonymous):

where do i plug them in and how do you figure that out lol

OpenStudy (anonymous):

I just want to know how to do the problem and then figure out which picture is correct and then finish this test its holding me up im so much work

OpenStudy (anonymous):

are you on the website? it should be the very first page there is y1 and y2 and it goes down from there. There are eye icons right beside where you would plug them in

OpenStudy (anonymous):

y less than or equal to x – 2 –4x + y more than sign – 3 you don't do anything with the Y just plug in x -2 beside the Y1= and -4x + y -3 in the Y2 = section and hit graph

OpenStudy (anonymous):

ok and how do i find the right picture ?

OpenStudy (anonymous):

graph actually, don't understand the graphs